OpenStudy (christos):

Find the line that contains the point P = (2,6) and is parallel to the line y = -4x + 26

hartnn (hartnn):

can you find slope of line y = -4x + 26 ?

OpenStudy (christos):

ok

hartnn (hartnn):

just compare it to y=mx+c , where m is the slope.

OpenStudy (christos):

-4

hartnn (hartnn):

yes.

hartnn (hartnn):

now parallel lines have same slope, so the slope of your line will also be -4, right ?

OpenStudy (christos):

hmm yea

OpenStudy (christos):

m will be same

hartnn (hartnn):

now u have a point and slope, can you find the equation ?

OpenStudy (christos):

I see lets try

hartnn (hartnn):

know the point slope form of equation of line ? or should i mention it ?

OpenStudy (skullpatrol):

Making a sketch may help.

OpenStudy (christos):

y=mx+b I replace y and x with 2 and 6 ? but then I would replace b too with 6 and there would be no equation, so I would do y=-4x+6 but that is still not the answer

OpenStudy (christos):

or it is indeed the answer ;p

OpenStudy (christos):

nah its not

hartnn (hartnn):

the form is y-y1 = m (x-x1) where point x1,y1 = 2,6 and m=-4

OpenStudy (christos):

isn't it y = mx + b ?

hartnn (hartnn):

thats slope intercept form :y=mx+b you have a point and slope , so, y-y1 = m (x-x1)

OpenStudy (christos):

I see

OpenStudy (skullpatrol):

"but then I would replace b too with 6" here is your mistake... you are solving for b in y=mx+b

OpenStudy (christos):

yea thats what I did indeed

OpenStudy (skullpatrol):

You do not replace b with 6

hartnn (hartnn):

because 'b' is y-intercept which is not given.

OpenStudy (christos):

I got y = mx - 2m +6

OpenStudy (skullpatrol):

You replace y with 6

hartnn (hartnn):

yes, just put m=-4 because you know it!

OpenStudy (christos):

:D

OpenStudy (christos):

y = -4x + 14 its correct :))

hartnn (hartnn):

:D good work!

OpenStudy (christos):

thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (christos):

now I have a similar one that is perpendicular this time instead of parallel

OpenStudy (christos):

I guess in that case m is not the same to work with that

hartnn (hartnn):

so you need to know that product of slopes of perpendicular lines = -1

hartnn (hartnn):

just to find slope.

OpenStudy (christos):

when you say "the product" of slopes you mean m1 + m2 = -1 ?

hartnn (hartnn):

product = multiplication :P m1 * m2 = -1

OpenStudy (christos):

aaah I see ok ^^

OpenStudy (whpalmer4):

It might be easier to think that the second slope will be -1/m if m is the slope of the first line...

OpenStudy (whpalmer4):

Completely equivalent to @hartnn's statement, but easier for some to grasp...

hartnn (hartnn):

yes^

OpenStudy (christos):

Nice one whpalmer sounds easier